Border Steel invested $500,000 in a new shearing unit. At an interest rate of 16% per year compounded quarterly, the annual income required to recover the investment in 3 years is closest to:

Question 11 options:

A)

$196,101

B)

$226,235

C)

$436,237

D)

$283,552
Answer #1

Effective interest = (1+0.16/4)^4 -1

= 1.04^4 -1

= 0.16985856

= 16.985856%

Annual income required = 500000 * (A/P, 16.985856%,3)

= 500000 * 0.452470

= 226234.98

= 226235

Option B is correct